Let X follows a normal distribution with mean 40 and variance 9. Then mean of Y= (X-40)/3 is


1
Expert's answer
2022-03-28T17:02:47-0400

"M({\frac {X-40} 3})=M({\frac X 3}-{\frac {40} 3})={\frac 1 3}M(X)-M({\frac {40} 3})={\frac {40} 3}-{\frac {40} 3}=0"